OpenStudy (anonymous):

what a^3-4a factored completley ?

OpenStudy (smurfy14):

a(a-2)(a+2)

OpenStudy (anonymous):

\[a^3-4a=a(a^2-4)=a(a+2)(a-2)\]

OpenStudy (anonymous):

thanks again LOl